The balance sheet for Tempest, Inc., is shown here in market value terms. There are 28,000 shares of stock outstanding. Markeshareholders equity by a. The transaction will b. New shares outstanding c. Share price
Answer #1

a). In lieu of a dividend of $1.60, the company has announced it is going to repurchase $44,800 worth of stock instead of paying a dividend.

It will reduce shareholders' equity by $44,800

b). P0 = Equity / Shares Outstanding

= $733,040 / 28,000 = $26.18

Shares Repurchased = Amount of Repurchasing the stock / P0

= $44,800 / $26.18 = 1,711.23

New shares outstanding = Original Shares Outstanding - Shares Repurchased

= 28,000 - 1,711.23 = 26,288.77

c). New Stock Price = [Original Equity - Repurchased Equity] / New Shares Outstanding

= [$733,040 - $44,800] / 26,288.77

= $688,240 / 26,288.77 = $26.18

> This is the right way to find your answer!

Justice7 Fri, Dec 3, 2021 7:38 AM